# Learning Dynamic Representations via An Optimally-Weighted Maximum Mean Discrepancy Optimization Framework for Continual Learning

### v4

@@ 1 Introduction
    We present an innovative framework termed Optimally-Weighted Maximum Mean Discrepancy (OWMMD) aimed at mitigating catastrophic forgetting in continual learning paradigms. The OWMMD framework utilizes a Multi-Level Feature Matching Mechanism (MLFMM) to impose penalties on the modification of feature representations across various tasks.
@@
    We introduce an Adaptive Regularization Optimization (ARO) framework that enables the model to evaluate the significance of each feature layer in real-time throughout the optimization process, thereby guaranteeing an optimal regularization process.
@@ 3.1 Problem definition
In a continual learning framework, the goal of the model is to identify the optimal parameter set ${\theta^{\star}}$ from the parameter space $\tilde{\Theta}$, which minimizes the training loss across all tasks ${T_{1},\cdots,T_{i}}$ while learning the current task $T_{i}$, as expressed by the following equation :
$\theta^{\star}=\underset{\theta\in\tilde{\Theta}}{\operatorname{argmin}}\frac{1}{i}\sum^{i}_{k=1}\Big{\{}\frac{1}{N^{s}_{i}}\sum^{N^{s}_{i}}_{j=1}\big{\{}{\mathcal{L}}\left({\bf y}^{i}_{j},f_{\theta}({\bf x}^{i}_{j})\right)\big{\}}\Big{\}},$ (1)
@@ 3.2 Multi-Level Feature Matching Mechanism
Let us formally define the function $F_{\theta_{i}}\colon{\mathcal{X}}\to{\mathcal{Y}}$ as a model parameterized by $\theta_{i}$, which comprises $K$ feature layers denoted as $\{F_{\theta_{i},1},F_{\theta_{i},2},\cdots,F_{\theta_{i},K}\}$, along with a linear classifier represented by $G_{\theta_{i}}\colon{\mathcal{Z}}\to{\mathcal{Y}}$. Here, the index $i$ signifies the model’s adaptation during the $i$-th task learning phase, and ${\mathcal{Z}}$ refers to the feature space corresponding to the output of the classifier. For a specified input ${\bf x}$, we can derive the feature vector from a particular feature layer through the following process :
$\mathcal{F}({\theta_{i}},{\bf x},k)=\begin{cases}F_{\theta_{i},1}({\bf x})&\text{if }k=1\\ F_{\theta_{i},k}\left(F_{\theta_{i},1}({\bf x})\right)&\text{if }k=2\\ F_{\theta_{i},k}\left(F_{\theta_{i},k-1}\left(\cdots F_{\theta_{i},1}({\bf x})\right)\right)&\text{if }3\leq k\leq K\,,\end{cases}$ (2)
where each feature layer $F_{\theta_{i},k}$ receives the output from the last layer $F_{\theta_{i},k-1}$ and produces the feature vector over the space ${\mathcal{Z}}^{k}$. The output of the classifier, or the logits, denoted as ${\bf z}\colon{\bf z}\in{\mathcal{Z}}$, is computed by applying the linear transformation $G_{\theta_{i}}$ to the final feature representation from the $K$-th layer:
${\bf z}=H_{\theta_{i}}(\mathbf{x})=G_{\theta_{i}}(F_{\theta_{i},K}(F_{\theta_{i},K}(\cdots F_{\theta_{i},1}({\bf x}))))\,,$ (3)
@@
${\bf\hat{y}}=F_{\theta_{i}}({\bf x})\triangleq softmax({\bf z})\,,$ (4)
where ${\bf\hat{y}}$ is the prediction for the input ${\bf x}$ and $C$ is the total number of classes. The $softmax()$ is the normalized function. For logits ${\bf z}=\{z_{1},\cdots,z_{C}\}$, we have :
$z_{t}=\frac{e^{z_{t}}}{\sum_{j=1}^{C}e^{z_{j}}},t=1,\cdots.C,$ (5)
In this paper, we adopt the cross-entropy loss function to update the parameter ${\theta_{i}}$, defined by :
@@ 3.3 The Maximum Mean Discrepancy based Regularization
Maximum Mean Discrepancy (MMD) serves as a significant and widely utilized distance metric within the realm of machine learning, grounded in a kernel-based statistical framework designed to assess the equivalence of two data distributions. Owing to its robust distance estimation capabilities, MMD has been adopted as a fundamental loss function for training diverse models across various applications, including image synthesis and density estimation. In contrast to alternative distance metrics such as Kullback–Leibler (KL) divergence and Earth Mover’s Distance, the principal advantage of the MMD criterion lies in its incorporation of the kernel trick, which facilitates the estimation of MMD on vectors without necessitating knowledge of the specific form of the density function.
The Maximum Mean Discrepancy (MMD) criterion is predicated on the concept of embedding probability distributions within a Reproducing Kernel Hilbert Space (RKHS). Let us denote $A$ and $B$ as two Borel probability measures. We introduce ${\bf a}$ and ${\bf b}$ as random variables defined over a topological space $\mathcal{X}$. Furthermore, we characterize the set $\{f\in\mathcal{F}|f\colon\mathcal{X}\to\mathbf{R}\}$ as a function, where $\mathcal{F}$ represents a specific class of functions. The MMD criterion quantifying the divergence between the distributions ${A}$ and ${B}$ is articulated as [57] :
${\mathcal{L}}_{\rm M}(A,B)\buildrel\Delta\over{=}\mathop{\sup}\limits_{f\in{\mathcal{F}}}\left({{{\mathbb{E}}_{{\bf a}\sim A}}\left[{f({\bf a})}\right]-{{\mathbb{E}}_{{\bf b}\sim B}}\left[{f({\bf b})}\right]}\right)\,.$ (8)
@@
where $\left\langle{f}\right.,\left.\cdot\right\rangle_{\mathcal{H}}$ is the inner product. Specifically due to the reproducing property of RKHS $f\in{\mathcal{F}},f({\bf a})=\left\langle{f}\right.,\left.{f_{k}({\bf a},\cdot)}\right\rangle_{\mathcal{H}}$, we can solve Eq. (9) by considering the kernel functions:
@@
Let $P_{{\bf Z}^{j}_{i}}$ and $P_{{\bf Z}^{j}_{i+1}}$ denote the distribution of ${\bf Z}^{j}_{i}$ and ${\bf Z}^{j}_{i+1}$, respectively. Based on the MMD criterion and Eq. (7), we can implement the regularization function when seeing the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$ at the new task learning $T_{i+1}$ as :
@@
${\mathbb{E}}_{({\bf x},{\bf y})\sim P_{D^{s}_{i+1}}}[{\mathcal{L}}_{s}({\bf y},F_{\theta_{i+1}}({\bf x}))]+\alpha\mathbb{E}_{({\bf x},{\bf z})\sim{P}_{{\mathcal{M}}_{i+1}}}[||{\bf z}-H_{\theta_{i+1}}({\bf x})||^{2}]$ (18)
@@
where $P_{D^{s}_{i+1}}$ and $P_{{\mathcal{M}}_{i+1}}$ denote the distribution of $D^{s}_{i+1}$ and the memory buffer ${\mathcal{M}}_{i+1}$ at the $(i+1)$-th task learning, ${\bf z}$ is the logits of previous task and was saved in the memory buffer. This approach helps the model retain past knowledge while learning new tasks efficiently.
@@
${\mathcal{L}}_{\rm total}$ $={\mathbb{E}}_{({\bf x},{\bf y})\sim P_{D^{s}_{i+1}}}[{\mathcal{L}}_{s}({\bf y},F_{\theta_{i+1}}({\bf x}))]+\alpha\mathbb{E}_{({\bf x},{\bf z})\sim{P}_{{\mathcal{M}}_{i+1}}}[||{\bf z}-H_{\theta_{i+1}}({\bf x})||^{2}]$ (19)
@@ 3.4 Adaptive Regularization Optimization Term
The final loss function is redefined as :
$\mathcal{L}^{\prime}_{\rm total}$ $={\mathbb{E}}_{({\bf x},{\bf y})\sim P_{D^{s}_{i+1}}}[\mathcal{L}_{s}({\bf y},F_{\theta_{i+1}}({\bf x}))]+\alpha\mathbb{E}_{({\bf x},{\bf z})\sim{P}_{{\mathcal{M}}_{i+1}}}[||{\bf z}-F_{\theta_{i+1}}({\bf x})||^{2}]$ (22)
@@ 3.5 Algorithm Implementation
1:  ${\bf Input:}$ dataset $D^{s}$, parameters $\theta$, scalars $\alpha$, $\beta$ and $\gamma$, learning rate $\eta$, number of model layers $K$
2:  ${\bf Init:}$${\mathcal{M}}\leftarrow\{\}$, ${\bf w}\leftarrow Uniform(0,1,K)$
3:  for $i=1$ to $N$ do
@@
9:        $\mathcal{L}_{\rm total}\leftarrow\mathcal{L}_{\rm total}+\alpha\cdot||{\bf z^{\prime}}-H_{\theta_{i}}({\bf x^{\prime}})||^{2}$
@@
6:  $d\leftarrow 0$
7:  ${\bf x^{\prime}}\leftarrow sample({\mathcal{M}})$
8:  $\mathcal{F}_{t}\leftarrow{\bf x^{\prime}}$
9:  $\mathcal{F}_{s}\leftarrow{\bf x^{\prime}}$

### v5

@@ 1 Introduction
    We propose a framework termed Optimally-Weighted Maximum Mean Discrepancy (OWMMD) aimed at mitigating catastrophic forgetting in continual learning paradigms. The OWMMD framework utilizes a Multi-Level Feature Matching Mechanism (MLFMM) to impose penalties on the modification of feature representations across various tasks.
@@
    To address over-regularization issues caused by the MLFMM, we propose to evaluate the significance of each feature layer when calculating the regularization loss term we call our approach as Adaptive Regularization Optimization (ARO). This approach can adaptively penalize significant changes on important feature layers while encouraging other feature layers to adapt to new tasks.
@@ 2 Related Work
Feature matching techniques have been introduced to mitigate catastrophic forgetting in continual learning, with a representative method outlined in [65]. This approach seeks to minimize the optimal transport distance between previous and current feature representations. However, computing optimal transport is often computationally prohibitive, and [65] addresses this challenge using the method from [1], which incurs significant computational overhead. In contrast, our work leverages Maximum Mean Discrepancy (MMD) for feature alignment, offering a more computationally efficient solution. Moreover, while [65] focuses solely on penalizing changes in the final feature outputs and overlooks intermediate representations, our proposed MLFMM framework enforces regularization across all feature layers, effectively alleviating network forgetting in continual learning scenarios. Additionally, we introduce an Adaptive Regularization Optimization (ARO) strategy that selectively constrains key feature layers, thereby preventing over-regularization. In contrast, [65] does not have such an adaptive feature matching mechanism and would easily lead to over-regularization issues.
@@ 3.1 Problem definition
In a continual learning framework, the goal of the model is to identify the optimal parameter set $\theta^{\star}$} from the parameter space $\tilde{\Theta}$, which minimizes the training loss across all tasks $\{{T_{1},\cdots,T_{i}}\}$ while learning the current task $T_{i}$, as expressed by the following equation :
$\theta^{\star}=\underset{\theta\in\tilde{\Theta}}{\operatorname{argmin}}\frac{1}{i}\sum^{i}_{k=1}\Big\{\frac{1}{N^{s}_{i}}\sum^{N^{s}_{i}}_{j=1}\big\{{\mathcal{L}}\left({\bf y}^{k}_{j},f_{\theta}({\bf x}^{k}_{j})\right)\big\}\Big\},$ (1)
@@ 3.2 Multi-Level Feature Matching Mechanism
Let us formally define the function $F_{\theta_{i}}\colon{\mathcal{X}}\to{\mathcal{Y}}$ as a model parameterized by $\theta_{i}$, which comprises $K$ feature layers denoted as $\{F_{\theta_{i},1},F_{\theta_{i},2},\cdots,F_{\theta_{i},K}\}$, along with a linear classifier represented by $G_{\theta_{i}}\colon{\mathcal{Z}}\to{\mathcal{Y}}$. Here, the index $i$ signifies the model’s adaptation during the $i$-th task learning phase, and $\mathcal{Z}$} refers to the feature space corresponding to the output of the feature layer. For a specified input $\bf x$}, we can derive the feature vector from a particular feature layer through the following process :
$\mathcal{F}({\theta_{i}},{\bf x},k)=F_{\theta_{i},k}\left(F_{\theta_{i},k-1}\left(\cdots F_{\theta_{i},1}({\bf x})\right)\right)\,$, (2)
where each feature layer $F_{\theta_{i},k}$ receives the output from the last layer $F_{\theta_{i},k-1}$ and produces the feature vector over the space ${\mathcal{Z}}^{k}$. The output of the final feature layer, denoted as ${\bf z}\colon{\bf z}\in{\mathcal{Z}}$, is computed by applying the linear transformations :
${\bf z}=F_{\theta_{i},K}(F_{\theta_{i},K-1}(\cdots F_{\theta_{i},1}({\bf x})))\,$, (3)
@@
${\bf\hat{y}}=Softmax(G_{\theta_{i}}({\bf z}))\,$, (4)
where $^{\bf\hat{y}}$ is the prediction for the input $\bf x$}. The $Softmax()$ is the normalized function. In this paper, we adopt the cross-entropy loss function to update the parameter $\theta_{i}$}, defined by :
@@
where $C$ represents the total number of categories. ${\bf y}[t]$ and ${\bf\hat{y}}[t]$ denote the $t$-th dimension of the real class label and the corresponding prediction.
@@ 3.3 The Maximum Mean Discrepancy based Regularization
Maximum Mean Discrepancy (MMD) [59] serves as a significant and widely utilized distance metric within the realm of machine learning, grounded in a kernel-based statistical framework designed to assess the equivalence of two data distributions. Owing to its robust distance estimation capabilities, MMD has been adopted as a fundamental loss function for training diverse models across various applications, including image synthesis and density estimation. In contrast to alternative distance metrics such as Kullback–Leibler (KL) divergence and Earth Mover’s Distance, the principal advantage of the MMD criterion lies in its incorporation of the kernel trick, which facilitates the estimation of MMD on vectors without necessitating knowledge of the specific form of the density function.
The Maximum Mean Discrepancy (MMD) criterion is predicated on the concept of embedding probability distributions within a Reproducing Kernel Hilbert Space (RKHS). Let us denote $A$ and $B$ as two Borel probability measures. We introduce $\bf a$} and $\bf b$} as random variables defined over a topological space $\mathcal{X}$. Furthermore, we characterize the set $\{f\in\mathcal{F}|f\colon\mathcal{X}\to\mathbf{R}\}$ as a function, where $\mathcal{F}$ represents a specific class of functions. The MMD criterion quantifying the divergence between the distributions A{A} and B{B} is articulated as [59] :
${\mathcal{L}}_{\rm M}(A,B)\mathrel{\mathop{\kern 0.0pt=}\limits^{\Delta}}\mathop{\sup}\limits_{f\in{\mathcal{F}}}\left({{{\mathbb{E}}_{{\bf a}\sim A}}\left[{f({\bf a})}\right]-{{\mathbb{E}}_{{\bf b}\sim B}}\left[{f({\bf b})}\right]}\right)\,.$ (7)
@@
where $\left\langle{f}\right$$\left.\cdot\right\rangle_{\mathcal{H}}$ is the inner product. Specifically due to the reproducing property of RKHS $f\in{\mathcal{F}},f({\bf a})=\left\langle{f}\right.,\left.{f_{k}({\bf a},\cdot)}\right\rangle_{\mathcal{H}}$, we can solve Eq. (8) by considering the kernel functions :
@@
Let $P_{{\bf Z}^{j}_{i}}$ and $P_{{\bf Z}^{j}_{i+1}}$ denote the distribution of ${\bf Z}^{j}_{i}$ and ${\bf Z}^{j}_{i+1}$, respectively. Based on the MMD criterion, we can implement the regularization function when seeing the data batch $\{{\bf x}_{1},\cdots,{\bf x}_{b}\}$ at the new task learning $T_{i+1}$ as :
@@
[…]
@@
where $P_{D^{s}_{i+1}}$ and $P_{{\mathcal{M}}_{i+1}}$ denote the distribution of $D^{s}_{i+1}$ and the memory buffer ${\mathcal{M}}_{i+1}$ at the $(i+1)$-th task learning, $\bf u$} is the logits of previous task and was saved in the memory buffer. This approach helps the model retain past knowledge while learning new tasks efficiently.
@@
[…]
@@ 3.4 Adaptive Regularization Optimization Term
By using the proposed AROT, the final loss function is redefined as :
[…]
@@ 3.5 Algorithm Implementation
[…]
3:for $i=1$ to NN do
@@
[…]
@@
[…]
@@
4 Theoretical Framework
In this section, we build upon the domain adaptation theory [41] to establish a novel theoretical framework for analyzing our proposed methodology. We begin by introducing key notations and definitions essential for the subsequent analysis.
Definition 1. Let us define $P_{{\bf Z}^{j}_{i}}$ and $P_{{\bf Z}^{j}_{i+1}}$ as the distribution of the feature vectors ${\bf Z}^{j}_{i}$ and ${\bf Z}^{j}_{i+1}$ produced by the $j$-th representation layer of the teacher and student model, respectively. Let $h\in{\mathcal{H}}$ denote a linear classifier and $\mathcal{H}$} as the space of classifiers.
Definition 2. Let $P_{{\bf Z}^{1:K}_{i}}$ denote the distribution of the unified representation that hierarchically integrates the feature vectors extracted by all feature layers. Based on $P_{{\bf Z}^{1:K}_{i}}$, the risk of the model can be defined by :
${\mathcal{E}}(P_{{\bf Z}^{1:K}_{i}},h,h^{\star})={\mathbb{E}_{{\bf z}\sim P_{{\bf Z}^{1:K}_{i}}}}[{\mathcal{L}}^{\prime}(h^{\star}({\bf z}),h({\bf z}))]\,$, (23)
where ${\mathcal{L}}^{\prime}(\cdot,\cdot)$ denotes a loss function that is symmetric and obeys the triangle inequality. $h^{\star}({\bf z})$ denotes a function that always returns the true class label for a given $\bf z$}.
Theorem 1. Let us define $h\in{\mathcal{H}}$ as a linear classifier. Let $P_{{\bf Z}^{1:K}_{i}}$ and $P_{{\bf Z}^{1:K}_{i+1}}$ denote two feature distributions formed using the teacher and student model, respectively. The loss function $\mathcal{L}$$^{\prime}$ is bounded ${\mathcal{L}}^{\prime}({\bf y}^{\prime},{\bf y}^{\prime})\leq M$, where $({\bf y},{\bf y}^{\prime})\in{\mathcal{Y}}^{2}$ and $M>0$ is a positive value. Let $h^{\star}_{{\bf Z}^{1:K}_{i}}=\arg\min_{h\in{\mathcal{H}}}{\mathcal{E}}(P_{{\bf Z}^{1:K}_{i}},h,h^{\star})$ and $h^{\star}_{{\bf Z}^{1:K}_{i+1}}=\arg\min_{h\in{\mathcal{H}}}{\mathcal{E}}(P_{{\bf Z}^{1:K}_{i+1}},h,h^{\star})$ denote the ideal classifiers for the feature distributions $P_{{\bf Z}^{1:K}_{i}}$ and $P_{{\bf Z}^{1:K}_{i+1}}$, respectively. We can derive a risk bound for the model at the $(i+1)$-th task learning, expressed as :
${\mathcal{E}}(P_{{\bf Z}^{1:K}_{i}},h,h^{\star})\leq{\mathcal{E}}(P_{{\bf Z}^{1:K}_{i+1}},h,h^{\star}_{{\bf Z}^{1:K}_{i+1}})+{\mathcal{D}}_{\rm disc}(P_{{\bf Z}^{1:K}_{i}},P_{{\bf Z}^{1:K}_{i+1}})+\epsilon\,$, (24)
where $\epsilon={\mathcal{E}}(P_{{\bf Z}^{1:K}_{i}},h^{\star},h^{\star}_{{\bf Z}^{1:K}_{i}})+{\mathcal{E}}(P_{{\bf Z}^{1:K}_{i+1}},h^{\star}_{{\bf Z}^{1:K}_{i+1}},h^{\star}_{{\bf Z}^{1:K}_{i}})$. ${\mathcal{D}}_{\rm disc}(P_{{\bf Z}^{1:K}_{i}},P_{{\bf Z}^{1:K}_{i+1}})$ is the discrepancy distance defined as :
$\mathcal{D}_{\rm disc}(P_{{\bf Z}^{1:K}_{i}},P_{{\bf Z}^{1:K}_{i+1}})$ $=\sup_{(h,h^{\prime})\in{\mathcal{H}}^{2}}\big|\mathbb{E}_{{\bf z}\sim P_{{\bf Z}^{1:K}_{i}}}[\mathcal{L}^{\prime}(h^{\prime}({\bf z}),h({\bf x}))]$ (25)
$-\mathbb{E}_{{\bf z}\sim P_{{\bf Z}^{1:K}_{i+1}}}[\mathcal{L}^{\prime}(h^{\prime}({\bf z}),h({\bf x}))]\big|\,.$
The proof is provided in [41]. In contrast to the risk bound described in [41], which solely assesses model performance on static target data distributions following training on static source data distributions, our proposed risk bound conceptualizes the preceding and current tasks as the source and target data distributions, respectively. This formulation enables a comprehensive analysis of forgetting phenomena within the context of continual learning. Specifically, Eq (24) quantifies the classifier’s performance on the preceding task (${T}_{i}$) following the update of the model parameters from ${\theta}_{i}$ to ${\theta}_{i+1}$ for the new task. According to Theorem 1, a substantial discrepancy term, ${\mathcal{D}}_{\rm disc}(P_{{\bf Z}^{1:K}_{i}},P_{{\bf Z}^{1:K}_{i+1}})$, elevates the left-hand side of Eq. (24), which represents the target model risk. This elevation signifies a decline in model stability. To mitigate this, we introduce a novel MLFMM designed to minimize the probabilistic divergence between the feature distributions of the previous and current tasks. This approach effectively reduces the discrepancy term ${\mathcal{D}}_{\rm disc}(P_{{\bf Z}^{1:K}_{i}},P_{{\bf Z}^{1:K}_{i+1}})$, thereby enhancing model stability. Nevertheless, excessive minimization of this term may constrain the model’s adaptability to novel tasks, impairing plasticity. To address this trade-off, we propose an AROT approach that selectively penalizes alterations in critical representation layers, thereby achieving a balance between plasticity and stability.
@@ 5.3 Comparison of Results for Complex Datasets
5.4 Compared to the Pre-Trained Models (PTMs)
Recent investigations have examined the pre-trained Vision Transformer (ViT) [15] as the core backbone architecture, incorporating a dynamically generated, lightweight expert module for learning new tasks. These pre-trained model (PTM)-based methodologies have demonstrated superior effectiveness in continual learning scenarios by leveraging the rich representations derived from PTMs. In this section, we extend the proposed OWMMD approach to PTM-based architectures with the objective of augmenting predictive accuracy and assessing the scalability of OWMMD. Specifically, consistent with [76], we utilize a pre-trained ViT as the foundational backbone and design an expert module comprising a feature extractor and a linear classifier. Upon encountering a novel task, a new expert module is instantiated dynamically, enabling focused learning of the task at hand, while all previously learned experts remain frozen to prevent catastrophic forgetting. To mitigate computational complexity, only the last three feature layers of the backbone are optimized during training. After completing each task, the primary backbone is duplicated and set to a frozen state, functioning as a teacher model. The OWMMD method is then applied to regulate the optimization process of the backbone’s parameters using the teacher module and the current active backbone. This modified approach, designated as OWMMD-PTMs, is evaluated against recent dynamic expansion models employing pre-trained ViT architectures, with detailed results presented in Table 6. Empirical findings indicate that OWMMD-PTMs consistently surpasses baseline methods across various datasets, substantiating its capacity to enhance model performance through the integration with pre-training strategies.
Table 6: The average accuracy calculated by various models on standard continual learning benchmarks, as averages of 10 runs. The results of baselines are taken from [6] and [64, 47, 76].
Methods
TinyImageNet CIFAR100
ICL w Pure-MM [47] - 96.35
KAMO [76] 92.05 $\pm$ 0.42 97.86 $\pm$ 0.37
DualPrompt [66] - 93.76
L2P [67] - 93.92
OWMMD-PTMs
94.15
$\pm$ 0.38
98.24
$\pm$ 0.32